Find the dimensions of the right circular cone of least volume that can be circumscribed about a sphere of radius a. Hint: x+2a=altitude of cone 


1
Expert's answer
2021-01-18T15:29:35-0500

In Triangle AOB,

"AO^2+AB^2=OB^2"

"(h-a)^2+r^2=a^2"

"\implies r^2 = a^2-(h-a)^2"


Volume of the Cone, "V = \frac{1}{3}\pi r^2 h = \frac{1}{3}\pi [ a^2-(h-a)^2 ]h"

For maximum and minimum,

"\frac{dV}{dh}=0"


Then,

"\frac{dV}{dh}= \frac{1}{3}\pi[ -2(h-a)h+a^2-(h-a)^2 ] =0"


"\implies h = \frac{4}{3}a"


Then, "r =\frac{2\sqrt{2}}{3}a"


Then, Volume will be, "V = \frac{32}{81}\pi a^3"